The 15-kg ball is connected with a spring as shown. Initially, the spring is compressed by 0.2m. the spring is stretched to 0.8 m. Determine the total work done to the ball during this process. the slope is smooth.
1
Expert's answer
2020-12-08T06:33:17-0500

Answer

Work done or stored energy in spring is given by

"W=E=-k\frac{x^2}{2}"

Here

Work done

"W=-k(\frac{(0.6)^2-(0.2)^2}{2}) =k(0.1)" J